2.0g mixture of MgCO3 and CaCO3 are heated till no further loss of weight takes place.The residue weighs 1.04g.Find the % composition of the mixture,(Mg=24,Ca=40,C=12,0=16)

Grade:12

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
total mass of mixture=2.0 g
let mass of MgCO3 = x g
So mass of CaCO3=2 -x g
decomposition reaction: MgCO3(s) -> MgO(s) + CO2 (g)
            84 g      40
CaCO3 (s) -> CaO(s) + CO2(g)
  100            56
so eqn,
84g MgCO3 leaves a residue =40 g
x g MgCO3 will leave residue = 40x/84 g
similarly for CaCO3 residue=56*(2-x) /100 g
Total = 40x/84 + 56*(2-x) /100 =1.04g -given
40 *100x + 84 *56 * 2 – 84 * 56x = 84 * 100 * 1.04
x= 0.96 g
%age of MgCO3= 0.96/2 *100 =48%
%age of CaCO3 = 100-48=52 % [Ans]
Hope it helped.
Thank you.
Soumya Ranjan Mohanty.